# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

### v2

@@ 6.3 Repeated Training Protocol
For each of the 10 dataset–model–split combinations (with five splits used for Random Forests), we perform independent training runs.
@@
varied deterministically over a log-spaced grid between and .
@@
Each training run produces a trained model, test accuracy, and a global explanation vector. Runs are executed in parallel on HPC infrastructure using chunked execution and aggregated post hoc (See the code repository). Due to the computational cost of Random Forest explanations, RF experiments were conducted over five independent train–test splits, which were sufficient to characterise the qualitative geometry of the explanation landscape.
Although the lbfgs solver is deterministic for fixed data and fixed hyperparameters, explanation multiplicity arises because training spans a range of regularisation strengths, allowing optimisation to settle into different equally optimal basins across the hyperparameter-conditioned loss landscape. Thus, the stochasticity measured by EvoXplain reflects basin selection under admissible training configurations, not randomness in optimisation or attribution.
@@ 6.5 Aggregation and Clustering
Clustering is performed using -means with candidate values
The optimal number of clusters is selected by maximising the silhouette score. Cluster centroids, sizes, and mechanistic entropy are computed as described in 4.3.
@@ 6.6 Disagreement Inspection
When , we identify representative runs for each explanatory basin by selecting the run closest to the cluster centroid in cosine distance.
Representative models are retrained and evaluated on the same test set. For each test instance , we compute the disagreement score
where denotes a representative model from basin .
@@ 6.7 Visualisation
For each split, explanation vectors are projected into two dimensions using PCA after centring. Per-split manifolds display cluster structure and regularisation gradients.
@@ 7 Results
7.1 Logistic Regression: explanation multiplicity under stochastic training
@@ 7.1.1 Breast Cancer
For a representative split (split seed 109), we analyse 1,000 independent training runs. In explanation space, the SHAP vectors do not form a single cloud. Instead, they organise along a smooth, low-dimensional arc with clear separation between regions (Fig. 1). Clustering in explanation space yields an optimal solution at , with a silhouette score of 0.776, indicating well-separated explanation modes. The normalised explanation entropy is 0.922, showing that runs are distributed across the two modes rather than collapsing into a single dominant explanation.
@@
The two clusters are not small perturbations of one another. Cluster 0 contains 663 runs with mean , while Cluster 1 contains 337 runs with mean . Despite occupying different regions of the explanation manifold, both clusters achieve comparable predictive accuracy (approximately 0.96). The accuracy distributions overlap substantially (Fig. 1, right panel). In other words, both explanation modes achieve similarly high predictive performance.
The clusters correspond to distinct regularisation regimes. The right panel of Fig. 1 shows a smooth gradient of values along the arc: strong regularisation (low ) concentrates at one end, while weak regularisation (high ) occupies the other. This gradient structure indicates that explanation geometry is tightly coupled to the regularisation–Loss landscape interaction.
These differences are reflected in the explanations themselves. The dominant features differ qualitatively between clusters. One mode emphasises classical size-related measurements, while the other places substantially more weight on variability- and shape-related features. The model is not merely reweighting the same features slightly differently; it is relying on qualitatively different explanatory patterns.
@@
Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—does not lie on the populated manifold. In the split-level embedding, it falls in a sparsely populated region between clusters (Fig. 1), indicating that this average explanation does not correspond to a typical trained model instance.
When explanation vectors are stacked across all ten splits, the same qualitative picture persists. The universal explanation manifold shows ten parallel arcs, each corresponding to a different train–test realisation, that remain separated in explanation space, while predictive accuracy remains uniformly high along all tracks (Fig. 2). The global average attribution vector, computed across all runs and all splits, lies again away from the high-density regions of the manifold. This reinforces that averaging explanations across stochastic runs produces a vector that is not representative of any deployed model instance.
@@
To assess whether these explanation differences translate into different predictions at the individual level, we compare representative models from the two clusters in split 109. In all held test instances, the mean absolute disagreement in the predicted malignant probability is 0.067 and the median is 0.035. However, the tail of the distribution is substantial: the maximum disagreement reaches , with 34 instances exceeding , 12 instances exceeding , and 4 instances exceeding . For the highest-disagreement case, one representative model assigns a malignant probability substantially different from the other, despite both models achieving high overall accuracy. These discrepancies arise even though aggregate performance metrics are indistinguishable.
@@
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength (). Across ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster (), with negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or dimensionality reduction, but emerges specifically when the training process admits multiple equally accurate models.
@@
Across splits, Logistic Regression on the Breast Cancer dataset consistently exhibits two well-separated explanation basins under stochastic training (Table 1). In every split, clustering selects explanation modes with high silhouette scores (0.76–0.82) and near-maximal normalised entropy (0.91–0.94), indicating that runs are distributed across multiple explanatory mechanisms rather than collapsing into a single dominant solution. This structure is stable across independent train–test realisations.
Importantly, predictive performance remains uniformly high across splits and clusters. Mean accuracy ranges from 0.956 to 0.976, with small standard deviations, and accuracy distributions overlap substantially between explanation basins within each split. There is no split in which explanation multiplicity coincides with degraded or unstable predictive performance, consistent with the split-level and universal explanation manifolds.
Cluster-level analysis (Table 2) further shows that the two explanation basins are not reducible to a simple partition by regularisation strength. Within each split, both clusters span the full sampled range of values, with similar mean and variance. This indicates that the mechanistic separation observed in explanation space reflects genuinely distinct explanatory modes rather than trivial hyperparameter binning. Note that while position along the arc correlates smoothly with regularisation strength (Fig.1, right panel), cluster membership does not partition by C: runs with identical regularisation can land in different basins depending on other aspects of the optimisation trajectory. As a control, when regularisation strength is fixed, explanations collapse to a single cluster across all splits, confirming that explanation multiplicity emerges specifically when the training process admits multiple equally accurate solutions.
@@ 7.1.2 COMPAS
Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4. Across all ten train–test splits, explanation clustering consistently selects two clusters (), with silhouette scores ranging from 0.73 to 0.91 (Table 3). These values indicate strong and stable separation between explanation modes across resampled datasets—notably higher than those observed on Breast Cancer. The normalised entropy ranges from 0.83 to 0.91, indicating that explanation modes are somewhat unequally populated, with cluster sizes typically in a ratio of approximately 1:2.5.
Predictive performance remains stable across splits. Mean accuracy ranges from 0.68 to 0.70 (this is typical for recidivism prediction tasks given inherent unpredictability), with standard deviations on the order of or smaller (Table 3). There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance. This is consistent with the split-level manifolds (Fig. 3), where accuracy varies smoothly along the explanation geometry rather than separating sharply by cluster.
Table 4 shows that, as with Breast Cancer, both explanation clusters span the full range of regularisation strengths. In every split, both clusters contain runs with values ranging from 0.01 to approximately 100, with similar mean values of around 10–12. This indicates that the mechanistic separation in explanation space is not a simple partition by regularisation strength. Rather, the two explanation modes represent genuinely distinct regions of the explanation landscape that models can reach regardless of their regularisation setting. Predictive accuracy remains effectively identical across clusters within each split, with differences typically below 0.001.
@@
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength (). Across all ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single cluster (), with zero entropy and negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or clustering procedure, but emerges specifically when the training process admits multiple equally accurate models through varied regularisation.
@@ 7.2 Disagreement Between Equally Accurate Explanation Basins
To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split. For each held-out test instance , we compute the absolute difference in predicted malignant probability,
@@
where and denote representative models selected from two different explanation clusters. Importantly, the compared models share the same model class, training data, and optimisation objective, and achieve comparable test accuracy. This controls for performance differences and isolates disagreement attributable to distinct explanatory mechanisms rather than optimisation noise or underfitting.
@@
On the Breast Cancer dataset, disagreement magnitudes span a wide range. While the mean disagreement across test instances is typically modest, the upper tail is substantial. Figure 5 illustrates this phenomenon for a representative split (Split 107), in which 76 test instances exhibit disagreement exceeding , 33 exceed , and 15 exceed . The maximum observed disagreement in this split reaches , despite both representative models achieving similar overall accuracy (approximately 0.95). These values indicate that explanation multiplicity can manifest as large probability divergences for individual patients even when aggregate performance metrics are indistinguishable.
@@
The highest-disagreement case in Figure 5 corresponds to a malignant tumour for which one model assigns a probability of 0.72 while the other assigns 0.04, leading to opposite clinical decisions for the same patient. Notably, this disagreement does not arise from marginal predictions near the decision threshold: one model is confidently malignant, while the other is confidently benign. The accompanying SHAP attributions reveal qualitatively different patterns of feature reliance across the two explanation basins. Although both models attend to overlapping morphological feature families, they weight and combine this evidence differently, resulting in incompatible diagnostic conclusions. This behaviour is inconsistent with an interpretation of disagreement as stochastic fluctuation or local instability, and instead reflects the presence of distinct explanatory mechanisms that remain viable at comparable accuracy.
@@
Across all ten train–test splits, the magnitude of individual-level disagreement varies substantially. Maximum disagreement per split ranges from negligible values (e.g., in split 103) to large discrepancies exceeding (e.g., splits 101 and 107). This variability indicates that the practical manifestation of explanation multiplicity depends on the particular data realisation. However, critically, large individual-level disagreements can arise even when aggregate accuracy and calibration metrics remain effectively unchanged. In such cases, standard performance summaries obscure the fact that different model instances may assign substantially different risk estimates to the same individual.
@@
Crucially, disagreement is not confined to cases where models predict different labels. Table 5 reports representative test instances from Split 107 for which both explanation basins correctly predict the same class, yet assign substantially different malignant probabilities (). In these cases, predictions agree at the label level, but differ meaningfully in confidence, with one model indicating borderline risk and the other indicating high certainty. While these instances do not lead to immediate label flips, they demonstrate that mechanistic disagreement can persist even when classification outcomes are identical, manifesting as divergent belief states rather than discrete decision changes. Such differences are relevant in downstream clinical contexts where predicted probabilities inform risk stratification, follow-up decisions, or human oversight.
@@
Taken together, these results show that explanation multiplicity is not merely a geometric or cosmetic property of explanation space. When multiple explanation basins coexist at similar accuracy, the choice of model instance can materially affect individual-level predictions, either by altering the predicted class or by substantially shifting estimated risk. The observed clusters therefore correspond to qualitatively distinct explanatory mechanisms, not small perturbations of a single underlying solution. High predictive performance alone does not imply mechanistic stability or explanatory uniqueness.

### v3

@@ 6.3 Repeated Training Protocol
For each dataset–model–split combination (10 splits for Logistic Regression and 5 splits used for Random Forests), we perform training runs.
@@
For Logistic Regression with varied , repeated runs correspond to repeated grid assignments (not stochastic optimisation), so variability reflects admissible pipeline variation along the regularisation path rather than run-to-run optimiser randomness.
@@
varied deterministically by assigning each run index to a value in a fixed log-spaced grid , cycling through as increases (so grid values repeat across the runs).
@@
Each training run produces a trained model, test accuracy, and a global explanation vector. Runs are executed in parallel on HPC infrastructure using chunked execution and aggregated post hoc (exploratory analyses in the repository). Due to the computational cost of Random Forest explanations, RF experiments were conducted over five independent train–test splits, which were sufficient to characterise the qualitative geometry of the explanation landscape.
Although the lbfgs solver is deterministic for fixed data and fixed hyperparameters, explanation multiplicity arises because the training pipeline spans a range of admissible regularisation strengths, each inducing a distinct optimal solution along the regularisation path. Thus, the variability measured by EvoXplain reflects selection among equally admissible solutions under the specified training configuration, rather than randomness in optimisation or attribution.
@@ 6.5 Aggregation and Clustering
We treat a single basin () as the null hypothesis corresponding to a unimodal explanation distribution.
For each candidate , we compute the silhouette score . A multi-basin solution is accepted only if , where we set as a conservative guardrail against over-partitioning. If no candidate meets this criterion, the explanation distribution is treated as unimodal and we set . In all experiments reported here, the minimum silhouette score among accepted multi-basin solutions exceeded 0.38, so conclusions do not hinge on the specific value of . To assess robustness to clustering choices, we additionally replicate basin detection using alternative clustering metrics and linkage criteria (Appendix A).
When one or more admissible multi-basin solutions exist, the optimal number of basins is selected by maximising the silhouette score among admissible . Basin centroids, sizes, and mechanistic entropy are then computed as described in Section 4.4.
While k-means assumes approximately spherical clusters, we use it here as a pragmatic heuristic for detecting multimodality in explanation space rather than as a geometric model of the manifold itself. In all reported cases, accepted multi-basin solutions exhibit high silhouette scores, indicating well-separated modes despite manifold curvature.
@@ 6.6 Disagreement Inspection
When , we identify representative runs for each explanatory basin by selecting the run closest to the basin centroid in cosine distance.
The corresponding representative trained models are evaluated on the same test set. For each test instance , we compute the disagreement score
where denotes a representative model instance drawn from basin .
@@ 6.7 Visualisation
For each split, explanation vectors are projected into two dimensions using PCA after centring. Per-split manifolds display basin structure and regularisation gradients.
@@ 7 Results
7.1 Logistic Regression: explanation multiplicity under admissible pipeline variation
@@ 7.1.1 Breast Cancer
For a representative split (split seed 109), we analyse 1,000 runs under the repeated training protocol. In explanation space, the SHAP vectors do not form a single cloud. Instead, for Logistic Regression they organise along a smooth, low-dimensional explanation manifold that traces a curved arc, within which recurrent regions of high density emerge with clear separation between them (Fig. 1). Clustering in explanation space yields an optimal solution at , with a silhouette score of 0.776, indicating well-separated explanation modes. The normalised explanation entropy is 0.922, showing that runs are distributed across the two modes rather than collapsing into a single dominant explanation.
@@
The two basins are not small perturbations of one another. Basin 0 contains 663 runs with mean , while Basin 1 contains 337 runs with mean . Despite occupying different regions of the explanation manifold, both basins achieve comparable predictive accuracy (approximately 0.96, See Table 1). In other words, both explanation modes achieve similarly high predictive performance.
Although regularisation strength varies smoothly along the explanation manifold (Fig. 1, right), the two explanation basins are not simple partitions of the sweep. Both basins span the full sampled range of regularisation values (Table 2), and runs with very similar can occupy different regions of the manifold. This overlap persists even when conditioning on narrow bands of regularisation strength, indicating that regularisation correlates with the continuous geometry of explanation space but does not uniquely determine basin membership. The mechanistic multiplicity therefore reflects non-identifiability under admissible training configurations, rather than trivial hyperparameter binning or noise.
These differences are reflected in the explanations themselves. The dominant features differ across basins: one mode places greater emphasis on classical size-related measurements, while the other assigns relatively more weight to variability and shape-related features. While these patterns arise within the same feature space, the resulting attribution profiles exhibit distinct explanatory emphases rather than simple uniform rescaling of a single feature ranking.
@@
Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—lies in a low-density region. In the split-level embedding, it falls in a sparsely populated region between basins (Fig. 2), indicating that this average explanation does not correspond to any high-density region of the explanation distribution.
When explanation vectors are stacked across all ten train–test splits and visualised jointly (without performing clustering), the same qualitative picture persists. In the resulting embedding, each split traces a distinct arc in explanation space corresponding to its specific train–test realisation (Fig. 1). These arcs remain separated from one another, even though predictive accuracy remains uniformly high along all trajectories.
@@
To assess whether these explanation differences translate into different predictions at the individual level, we compare representative models from the two basins in split 107. While the mean absolute disagreement in predicted malignant probability across test instances is modest, the upper tail of the distribution is substantial. In this split, 76 test instances exhibit disagreement exceeding , 33 exceed , and 15 exceed , with the maximum observed disagreement reaching . For the highest-disagreement case, one representative model assigns a malignant probability of 0.72 while the other assigns 0.04, leading to opposite predicted labels under a 0.5 threshold, despite both models achieving similarly high overall accuracy. These discrepancies arise even though aggregate performance metrics are effectively indistinguishable.
@@
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength (). Across ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single basin (), with negligible dispersion in explanation space. This indicates that, in this setting, explanation multiplicity arises under admissible hyperparameter-conditioned training configurations (varied ), rather than from randomness in the optimiser or attribution method. (This is true for Adult Income dataset as well, which can be found in Appendix A)
Orthogonal Synthetic Control. To rule out feature correlation as a driver of explanation multiplicity, we repeat the analysis on a synthetic dataset with strictly orthogonal features and a known linear ground-truth rule. Under fixed regularisation, explanations collapse to a single mode (). When regularisation is allowed to vary across admissible values, explanations again organise into a structured low-dimensional manifold with recurrent regions, despite the absence of correlated or redundant features. This demonstrates that explanation multiplicity does not rely on feature correlation or dataset-specific quirks, but can arise even in idealised settings where explanatory ambiguity would not ordinarily be expected. Full details are provided in Appendix A.
@@
A potential alternative explanation is that the observed basins merely reflect different regularisation regimes. However, this is not supported by the data. As shown in Table 2, the two explanation basins identified for Split 100 have nearly identical mean regularisation strengths (), yet occupy well-separated regions of explanation space. This demonstrates that explanation multiplicity is not reducible to a monotonic effect of regularisation strength, but instead reflects distinct explanation modes compatible with similar model configurations.
Across splits, Logistic Regression on the Breast Cancer dataset consistently exhibits two well-separated explanation basins under admissible training pipeline variation (Table 1). In every split, clustering selects explanation modes with high silhouette scores (0.76–0.82) and near-maximal normalised entropy (0.91–0.94), indicating that runs are distributed across multiple explanation basins rather than collapsing into a single dominant pattern. This structure is stable across independent train–test realisations.
Importantly, predictive performance remains uniformly high across splits and basins. Mean accuracy ranges from 0.956 to 0.976, with small standard deviations, and accuracy distributions overlap substantially between explanation basins within each split. There is no split in which explanation multiplicity coincides with degraded or unstable predictive performance, consistent with the split-level and universal explanation manifolds.
Basin-level analysis (Table 2) further shows that the two explanation basins are not reducible to a simple partition by regularisation strength. Within each split, both basins span the full sampled range of values, with similar mean and variance. This indicates that the mechanistic separation observed in explanation space reflects genuinely distinct explanatory modes rather than trivial hyperparameter binning. As a control, when regularisation strength is fixed, explanations collapse to a single basin across all splits, confirming that explanation multiplicity emerges specifically when the training process admits multiple equally accurate solutions.
@@ 7.1.2 COMPAS
Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4. Across all ten train–test splits, explanation clustering consistently selects two basins (), with silhouette scores ranging from 0.73 to 0.91 (Table 3). These values indicate strong and stable separation between explanation modes across resampled datasets—notably higher than those observed on Breast Cancer. The normalised entropy ranges from 0.83 to 0.91, indicating that explanation modes are somewhat unequally populated, with basin sizes typically in a ratio of approximately 1:2.5.
Predictive performance remains stable across splits. Mean accuracy ranges from 0.68 to 0.70 (consistent with typical reported performance ranges for this task), with standard deviations on the order of or smaller (Table 3). There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance. This is consistent with the split-level manifolds (Fig. 3), where accuracy remains confined to a narrow range across the explanation geometry rather than separating sharply by basin.
Table 4 shows that, as with Breast Cancer, both explanation basins span the full range of regularisation strengths. In every split, both basins contain runs with values ranging from 0.01 to approximately 100, with similar mean values of around 10–12. This indicates that the mechanistic separation in explanation space is not a simple partition by regularisation strength. Rather, the two explanation modes represent genuinely distinct regions of the explanation landscape that models can reach regardless of their regularisation setting. Predictive accuracy remains effectively identical across basins within each split, with differences typically below 0.001.
@@
Control Experiment: Stability under Fixed Regularisation. As a control, we repeat the same analysis under fixed regularisation strength (). Across all ten independent train–test splits, each with 1,000 runs, the explanation vectors consistently collapse to a single basin (), with zero entropy and negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or clustering procedure, but emerges specifically when the training process admits multiple equally accurate models through varied regularisation.
@@ 7.2 Disagreement Between Equally Accurate Explanation Basins
To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split. For each test instance , we compute the disagreement score
where denotes a representative model selected from explanatory basin . In the present experiments, clustering yields basins, so this reduces to
@@
Importantly, the compared models share the same model class, training data, and optimisation objective, and achieve comparable test accuracy. This controls for trivial performance differences, ensuring that observed disagreement reflects differences in admissible explanatory behaviour under the training pipeline rather than underfitting or attribution noise.
@@
On the Breast Cancer dataset, disagreement magnitudes span a wide range. While the mean disagreement across test instances is typically modest, the upper tail is substantial. Figure 5 illustrates this phenomenon for a representative split (Split 107), in which 76 test instances exhibit disagreement exceeding , 33 exceed , and 15 exceed . The maximum observed disagreement in this split reaches , despite both representative models achieving similar overall accuracy (approximately 0.95). These values indicate that explanation multiplicity can manifest as large probability divergences for individual patients even when aggregate performance metrics are indistinguishable. Split 107 is shown as an illustrative high-disagreement example; other splits exhibit similar behaviour with varying magnitude.
@@
The highest-disagreement case in Figure 5 corresponds to a malignant tumour for which one model assigns a probability of 0.72 while the other assigns 0.04, leading to opposite predicted labels under a 0.5 threshold for the same patient. Notably, this disagreement does not arise from marginal predictions near the decision threshold: one model is confidently malignant, while the other is confidently benign. The accompanying SHAP attributions reveal qualitatively different patterns of feature reliance across the two explanation basins. Although both models attend to overlapping morphological feature families, they weight and combine this evidence differently, resulting in incompatible diagnostic conclusions. This pattern persists across multiple representative pairs within the same basins (exploratory analyses in the repository).
@@
Across all ten train–test splits, the magnitude of individual-level disagreement varies substantially. Maximum disagreement per split ranges from negligible values (e.g., in split 103) to large discrepancies exceeding (e.g., splits 101 and 107). This variability indicates that the practical manifestation of explanation multiplicity depends on the particular data realisation. However, critically, large individual-level disagreements can arise even when aggregate accuracy and aggregate performance metrics remain effectively unchanged. In such cases, standard performance summaries obscure the fact that different model instances may assign substantially different risk estimates to the same individual.
@@
Crucially, disagreement is not confined to cases where models predict different labels. Table 5 reports representative test instances from Split 107 for which both explanation basins correctly predict the same class, yet assign substantially different malignant probabilities (–). In these cases, predictions agree at the label level, but differ meaningfully in confidence, with one model indicating borderline risk and the other indicating high certainty. While these instances do not lead to immediate label flips, they demonstrate that explanation-level disagreement can persist even when classification outcomes are identical, manifesting as divergent confidence assessments rather than discrete decision changes. Such differences are relevant in downstream clinical contexts where predicted probabilities inform risk stratification, follow-up decisions, or human oversight.
@@
Taken together, these results show that explanation multiplicity is not merely a geometric or cosmetic property of explanation space. When multiple explanation basins coexist at similar accuracy, the choice of model instance can materially affect individual-level predictions, either by altering the predicted class or by substantially shifting estimated risk. The observed basins therefore correspond to qualitatively distinct explanatory mechanisms, not small perturbations of a single underlying solution. High predictive performance alone does not imply mechanistic stability or explanatory uniqueness.